a weight is suspended from a steel wire of length 1.0m and cross sectional area 1.0mm². The ultimate strength of the steel is 400.0 x 10^6 N.m². What is the maximum mass that the weight can have without breaking the wire?
1
Expert's answer
2019-12-04T08:40:43-0500

The ultimate strength of the steel is(P)= 400.0 x 106 N/m²

Cross section area(A)= 1mm2=1/10002m2

Thus total strength(F)is

"F=P\times A=4000\times10^6\times1000^{-2}=400N" N

"F=m\times g=m\times9.81"

"400=m\times 9.81"

m=40.77kg